OpenStudy (anonymous):

A car is traveling 80.1 kilometers per hour. What is its speed in miles per minute? One kilometer = 0.62 miles.

OpenStudy (anonymous):

80.1 x 0.62 / 60 = 0.8277miles per min

OpenStudy (anonymous):

80.1 is in kph, not mpm

OpenStudy (anonymous):

80.1 km per hour = 80.1/60km per min = 1.335 km per min 0.62 miles = 1km so 1.335 x 0.62 = 0.8277miles per min

OpenStudy (anonymous):

\[\frac{80.1kilometers}{1hour}*\frac{.062 miles}{1 kilometer}*\frac{1 hour}{60 minutes}\]

OpenStudy (anonymous):

kilos cancelt hours cancel you're left with miles /mins